Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?

Answer by ewatrrr(24785)


Hi,
Let x represent the time in minutes it takes all three to fill the pool
Per minute is the Equalizer***
1/30 + 1/45 + 1/90 = 1/x Multiplying all terms by 90x so as all denominators= 1
3x + 2x + x = 90
6x = 90
x = 15 minutes, the time it takes all three to fill the pool
chaecking our answer
3/90 + 2/90 + 1/90 = 6/90 = 1/15